Massachusetts Institute of Technology Handout 1 Mathematics for Computer Science August 5, 2002 Prof. Charles E. Leiserson

Minicourse Information

MIT Subject

This minicourse is based on the MIT subject 6.042J/18.062J Mathematics for Computer Science, which covers discrete mathematics and probability. The material in the MIT subject 6.042 is a prerequisite for SMA5503.

Staff

MIT Instructor Charles E. LEISERSON cel@mit.edu
NUS Instructor LEE Wee Sun  leews@comp.nus.edu.sg

        TA  Kunal AGRAWAL  smap1071@nus.edu.sg

Course Secretary Leigh DEACON  deacon@mit.edu

Schedule

Lectures will be held in room 34-101, not the regular classroom.

Proofs, Induction  Monday, August 5 10:30 – noon
Probability  Tuesday, August 6 10:30 – noon
Probability, Independence Thursday, August 8 10:30 – noon
Expectation  Friday, August 9  10:30 – noon

How to Learn More

Two textbooks are used for the full-length version of this course at MIT. The book The Nuts and Bolts of Proofs by Antonella Cupillari is a small paperback with content similar to our first lecture. The book Discrete Mathematics and Its Applications by Kenneth Rosen is an 800-page text. Sections 3.1 and 3.2 of Rosen’s book cover our first lecture, and sections 4.4 and 4.5 (very briefly) cover the second and third.

Written notes from the various 6.042 course offerings are available at:

             http://theory.lcs.mit.edu/classes/6.042/spring02/

Particularly relevant are Weeks 1 and 2 (on proofs and induction), Weeks 9 and 10 (on probability), and Week 11 (on expectation). You must know all the material in 6.042 as a prerequisite for SMA5503, even though we will be reviewing only part of the material.